a sample of 91 circuits from a large normal population has a mean resistance of 2,4 ohms. if the population standard deviation is 1.5 ohms, determine a 95% confidence interval for the true mean resistance of the population. (hint: critical value is 1.96) what is the lower limit value


1
Expert's answer
2021-07-05T18:10:12-0400

The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times \dfrac{\sigma}{\sqrt{n}},\bar{x}-z_c\times \dfrac{\sigma}{\sqrt{n}})"

"=(2.4-1.6\times \dfrac{1.5}{\sqrt{91}}, 2.4+1.6\times \dfrac{1.5}{\sqrt{91}})"

"=(2.0918, 2.7082)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "2.0918<\mu<2.7082," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(2.0918, 2.7082)."